40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 17 15:17:16 2016, gautam, Update, IMC, MCL Feedback MCLerror.pdfMCLtest.pngYarmCtrl.pdf
    Reply  Wed Nov 23 01:13:02 2016, gautam, Update, IMC, MCL Feedback MCL_plant.pdfOLG.pdfMC_armSpectra_X.pdfMC_armSpectra_Y.pdf
       Reply  Wed Nov 23 15:08:56 2016, rana, Update, IMC, MCL Feedback 
          Reply  Fri Feb 3 00:36:52 2017, gautam, Update, IMC, MCL Feedback - framing the problem 
             Reply  Fri Feb 3 11:40:09 2017, rana, Update, IMC, MCL Feedback - framing the problem 
                Reply  Mon Feb 6 17:03:41 2017, gautam, Update, IMC, MCL Feedback - simulink model updated mc40_v1.pdf
                   Reply  Mon Feb 6 18:20:08 2017, Koji, Update, IMC, MCL Feedback - simulink model updated 
                      Reply  Wed Feb 8 19:13:02 2017, gautam, Update, IMC, MCL Feedback - TF measurements 
                         Reply  Thu Feb 9 23:35:34 2017, gautam, Update, IMC, MCL Feedback - TF measurements mc40_v1.pdfCMboard_OLTF_fit.pdfFSSFast_OLTF_fit.pdfPCdrive_OLTF_measured.pdfdata.zip
Message ID: 12812     Entry time: Wed Feb 8 19:13:02 2017     In reply to: 12805     Reply to this: 12815
Author: gautam 
Type: Update 
Category: IMC 
Subject: MCL Feedback - TF measurements 

Quick summary elog, details to follow. I did the following:

  • Updated the Simulink model based on Koji's feedback. 
  • Today morning, I measured the (electronic) open-loop TFs of
    • MC Servo Board
    • FSS Fast path (PZT)
    • FSS PC Drive path
  • The summing amplifiers in the latter two paths are assumed to be broadband for the purposes of this model.

The measurements I have look reasonable. But I had a hard time trying to look at the schematic and determine what is the appropriate number and locations of poles/zeros with which to fit the measured transfer function. Koji and I spent some time trying to go through the MC Servo board schematic, but looks like the version uploaded on the 40m DCC tree doesn't have changes made to it reflected (we compared to pictures on the 40m google photos page and saw a number of component values were different). Since the deviation between fit and measurement only occurs above 1MHz (while using poles/zeros inferred from the schematic), we decided against pulling out the servo board and investigating further - but this should be done at the next opportunity. I've marked the changes we caught on a schematic and will upload it to the 40m DCC page, and we can update this when we get the chance.

So it remains to fit the other two measured TFs, and add them to the Simulink model. Then the only unknown will be the PDH discriminant, which we anyway want to characterize given that we will soon have much more modulation.  

Data + plots + fits + updated schematics to follow...

 

ELOG V3.1.3-